# Does Positive Reinforcement Work?: A Quasi-Experimental Study of the Effects of Positive Feedback on Reddit

### v1

@@ 4.1.1. Findings
Table 1. Difference-in-differences regression analysis results. We report the percent increase observed in the dependent variable for each treated user compared to if they had never received the treatment. Regressions with Bonferroni-corrected $p$-values are reported with asterisks (* $p<0.05$, ** $p<0.01$, *** $p<0.001$). Cohen’s $d$ reports the effect size of treatment on the dependent variables. The table shows significant increases in score and gilds as a result of either treatments, and both an increase in daily post rate and a decrease in daily removal rate as a result of the score treatment.
@@
Gild Avg. Score 19.28% *** 0.16
Avg. # Gilds 1.63% *** 0.27
@@
The results of our DiD analysis are reported in Table 1. The increase column is calculated from the coefficient on the $a_{t}*\delta(d\geq 0)$ term which indicates the effect of treatment on treated users in the observation window. To reduce the likelihood of spurious significant relationships, we perform a Bonferroni correction by multiplying the $p$-values of each coefficient by the product of the number of regression tests (6) and the number of independent variables per regression (4). To capture the effect size, we also report the Cohen’s $d$ measured between the difference in each dependent variable before and after treatment for each treated user compared to the same difference before and after placebo for each control user.
We find that both forms of treatment lead to large positive increases in treated users’ score on future posts compared to if they had not received the treatment. More specifically, receiving a gild and being highly upvoted correlate with a user receiving 19.28% and 57.11% higher score than they would have, respectively. Both treatments have much more modest, but still significant, effects on the average number of gilds received on future posts. The gild treatment encourages 1.63% more gilds in the future while the score treatment encourages an increase of less than 1% more gilds.
Users treated with score post 2.24% more per day than they would have if they had not been treated. Additionally, receiving high score on a post has a negative effect on users’ removal rate. This demonstrates that receiving a lot of upvotes not only encourages more positive contributions, but also decreases the rate of norm-violating behavior.
@@
Overall, we see that receiving positive feedback of either kind encourages users to contribute posts that are better-received by their communities, and in some cases, also promotes user engagement and norm acquisition. Given the increase in future positive feedback and the decrease in norm-violations, we say that being highly upvoted encourages users to post higher-quality content in the eyes of both the community and the centralized moderation team.
@@ 4.2. RQ1b: How long do the effects of positive feedback on user behavior last?
Figure 5. Average treatment effects (ATE) of each treatment on the four outcomes. The rightmost $y$-axes correspond to the ATE of the score outcome which has a much larger range. The “x” markers indicate the days at which each outcome measure reaches saturation (i.e., decreases or stays the same for two consecutive days). The figure demonstrates the lasting effect of the score treatment on each outcome compared to the gild treatment, where the ATEs mostly return to zero by the end of our observation window.
@@
A plot with two subplots. Both plots show the average treatment effects for each outcome from the start of the baseline period to the end of the observation window. The leftmost plot shows these values for the gild treatment and the rightmost plot visualizes the daily ATE values for the score treatment. The ’x’ icons on the leftmost plot show that the ATE for daily removal rate reaches saturation first, followed by daily post rate, average score, and finally average number of gilds. In the rightmost plot, removal rate reaches saturation first, then average score, daily post rate, and average number of gilds.
To determine how long the effect of each treatment lasts (RQ1b), we compute the Average Treatment Effect (ATE) for each of our four outcomes on each day in our study period. ATE is computed as the difference between the treated users’ change in outcome and the control users’ change in outcome before and after treatment. We visualize the ATE values in Figure 5 averaged across each day in the baseline and observation windows. For each outcome, we also identify the day at which it reaches saturation, meaning the ATE for the outcome stops increasing for two consecutive days after treatment. Formally, our saturation point is defined as the day $d$ such that $ATE(d)\leq ATE(d-1)\leq ATE(d-2)$. For the removal rate outcome, we reverse the definition of saturation point since we are interested in seeing decreases in norm violations. In other words, the saturation point is the first day $d$ for which $ATE(d)\geq ATE(d-1)\geq ATE(d-2)$. These points are denoted by red “x” symbols in Figure 5.
We observe that for the gild treatment, the ATE values for all four outcomes return to nearly zero by the end of the observation window, indicating minimal differences between treated and control users. For the users treated with score, however, the ATE of their score and post rate in the observation window remain well above zero at the end of the two-week observation window, even though both measures reach their saturation point by day 5. This indicates that the effect of being highly-upvoted on future upvotes and engagement may last longer than the effect of being gilded. The elevated scores in the observation window may also serve as additional treatments, creating a snowball effect.
@@ 4.2.1. Rebound
We notice that most treated users do not rebound on the score outcome within 14 days, reflecting the trend shown in Figure 5 that score is elevated throughout the entire observation window. This trend is also observed for daily post rate regardless of the treatment. However, for the average number of gilds, nearly all users in both treatments rebound to their baseline values within the two-week window. This is understandable given the rarity of gilding in the dataset.
Finally, we observe that the percentage of users who rebound to baseline removal rates after either treatment is much more evenly split. This suggests that the treatments teach some users about the norms more effectively than others. Our subsequent RQ2 analysis explores this trend in more depth.
Table 2. Summary table describing the number of treated users that rebound. By definition, users must post at least twice in the 2-week window after their treatment in order to rebound, which is true for 2,599 (60.2%) of users treated with gilds and 90,993 (35.27%) users treated with high score. The table highlights the fact that users almost always rebound on their average number of gilds, rarely rebound on their average score and daily post rate, and rebound on their daily removal rate roughly half the time.
@@
Gild Avg. Score 74 (1.72%) 2,525 (58.53%)
Avg. # Gilds 2,547 (59.04%) 52 (1.21%)
@@
5. RQ2: How do the effects of positive feedback vary between recipients?
@@ 5.1. RQ2a: What types of users experience stronger effects of receiving positive feedback?
We compute Individual Treatment Effects (ITE) of each of our treatments on all four outcomes to evaluate user-level differential effects. ITE is used in prior work to compute how much a single treated user is affected by the treatment (Saha et al., 2019b). In our work specifically, ITE is computed as the difference between a treated user’s change in behavior after treatment and the average change in all control users’ behavior from the same stratum after placebo. More precisely, for an individual treated user from stratum $s$, $u_{t,s}$, we compute the ITE for each outcome as follows:
$\textit{ITE}_{\text{outcome}}(u_{t,s})=\Delta u_{t,s}(\text{outcome})-\text{mean}(\Delta u_{c,s}(\text{outcome}):u_{c,s}\in\mathcal{D}_{\textit{control,}s})$
@@
We then perform one linear regression analysis for each treatment-outcome pair using the outcome’s ITE as the dependent variable. Baseline values for our four outcomes are independent variables along with the author’s age, karma (i.e., the sum of scores on all posts) in the month prior to the baseline window, stratum, and the strength of treatment (i.e., number of gilds or score on treated post). We take the log of each numerical independent variable.
@@ 5.1.1. Findings
Table 3 reports the coefficients from this regression. The bold coefficients are those we discuss in this section. All significance tests are Bonferroni-corrected based on the number of independent variables and regressions run. For the gild treatment, each $p$-value is multiplied by 28 (4 regressions, 7 IVs) and for score, each $p$-value is multiplied by 32 (4 regressions, 8 IVs). This regression analysis demonstrates that the strength of either treatment may lead to stronger effects. For example, posts treated with more gilds are positively correlated with the treated user experiencing stronger effects on their future number of gilds. Similarly, users who are treated with higher score have more extreme effects on their future positive feedback than users with more modest treatments. With this score treatment, however, we see that being treated with lower score also correlates with a larger effect on removal rate.
We also see trends aligned with the ideas of satiation and deprivation introduced by Miltenberger (2016). Specifically, our regression shows that users who have higher baseline average score experienced diminished effects of either treatment on their future score. Similarly, users with a higher average number of gilds in their baseline window see less dramatic effects of the score treatment on their future number of gilds. Both these trends reflect the idea that users are more strongly impacted by stimuli (e.g., gilds or upvotes) if they have been deprived of that stimulus.
@@
Table 3. Individual Treatment Effects (ITE) regression results. Cells that are grayed out indicate independent variables not present in the regression. Regressions with Bonferroni-corrected $p$-values (* $p<0.05$, ** $p<0.01$, *** $p<0.001$). Coefficients with $p>0.05$ are excluded. Bold coefficients are discussed in Section 5.1.1.
Treatment $\rightarrow$ Gild Score
IVs $\downarrow$ ITE $\rightarrow$ Score # Gilds Post Rate Rem. Rate Score # Gilds Post Rate Rem. Rate
Author Age -9.7E-03 *** -2.4E-03 ***
@@
Baseline Post Rate -2.3E-02 *** -8.8E-01 *** -4.6E-03 *** -2.0E-01 *** 3.3E-01 ***
Baseline Removal Rate -7.1E-02 * -2.6E-01 *** 2.3E+00 *** -3.1E-01 *** -2.5E+00 ***
Karma Last Month 1.1E-01 *** 1.8E-02 *** 9.1E-04 *
Gilds Received 3.0E-02 **
Score Received 9.7E-02 *** 5.6E-04 *** -2.6E-03 ***
Baseline Avg. # Gilds -5.0E-01 ***
R-squared 0.110 0.052 0.109 0.076 0.108 0.182 0.057 0.244
@@
Regression results for the eight regressions measuring the effect of the independent variables on the ITE values for each treatment-outcome pair.
@@ 5.2. RQ2b: What types of communities experience stronger average effects of receiving positive feedback?
$\textit{CTE}_{\text{outcome}}(r)=\text{mean}(\Delta u_{t,r}(\text{outcome}):u_{t,r}\in\mathcal{D}_{\textit{treated,}r})-\text{mean}(\Delta u_{c,r}(\text{outcome}):u_{c,r}\in\mathcal{D}_{\textit{control,}r})$
@@
Similar to the previous ITE analysis, we use our CTE values as dependent variables for eight linear regression analyses with subreddit-level monthly activity statistics from our entire study period as independent variables. This includes post rate, average number of gilds, average score, the number of unique authors, and the number of subscribers.
@@ 5.2.1. Findings
After Bonferroni correction, these regressions yielded only one significant coefficient revealing that subreddits that gild more often see larger effects of the score treatment on the their future number of gilds. The lack of significant results indicates that the treatments have similar effects regardless of the community in which they occur.
@@ 6. Discussion
Our work contributes a computational approach and a causal inference framework through which researchers can examine the differential effects of treatment outcomes. We outline a method of identifying treatment and control posts, performing a difference-in-differences analysis, and determining user- and community-level factors that influence the strength of the treatment effect. This methodology can be adapted with different treatments, outcomes, and platforms to fit researchers’ needs. Additionally, while we focused on two types of positive feedback available on Reddit, upvotes in particular are almost universally applicable across social media platforms and have analogous features on Facebook, Instagram, X, Stack Overflow, and YouTube, among others. This ubiquity of upvotes enhances the generalizability of our work. Researchers interested in other platforms can use these findings to inform research into analogous features and extend this work with analyses of positive feedback signals not available on Reddit.
@@
6.1. Prioritize highlighting content contributed by users recently deprived of positive feedback.
A key differential effect we observed relates to the principle of deprivation in positive reinforcement theory. The principle states that a stimulus has larger impacts on subjects deprived of it (Miltenberger, 2016). Our findings confirm this theory in the context of positive feedback on Reddit, highlighting that users who had not recently received a gild experienced stronger effects of being gilded on their future number of gilds. Similarly, users who had not recently been highly-upvoted experienced larger increases in their future score when experiencing the score treatment.
@@
We recommend that designers and HCI researchers consider the principles of satiation and deprivation to best utilize valuable community resources like time and effort. Specifically, we suggest that they consider the impact of deprivation when highlighting content in social media feeds to maximize the potential effects of positive feedback, especially since both moderators and users have limited time and effort to dedicate to a community.
6.2. Provide mechanisms to bring focused attention on high-quality newcomer contributions.
@@
We showed that positive feedback in the form of upvotes results in the recipient posting fewer norm-violating posts than they would have without the treatment, implying that recipients are learning community norms. Additionally, newer accounts see stronger effects on their norm-adherence after treatment compared to more established Reddit users indicating that platform-level newcomers benefit most from positive feedback in the context of norm acquisition.
@@
To enable this norm acquisition, we call on designers and HCI researchers to explore explicit ways for moderation teams to encourage established community-members to review newcomer contributions and provide positive feedback when applicable. This may be a highlighting mechanism, a pop-up window nudging users, or a dedicated feed of newcomer contributions within communities.
@@
User motivation can be intrinsic or extrinsic (Kraut et al., 2011). In the context of our research, intrinsically-motivated users are those who contribute high-quality content because they enjoy doing so. Kraut et al. (2011) highlight performance-based positive feedback as an especially effective method of enhancing users’ intrinsic motivation. Therefore, users who receive community-wide approval through positive feedback may experience increased intrinsic motivation to participate.
Extrinsically-motivated users, may be more interested in making high-quality posts as a means to increase their reputation (i.e., karma) in the community. The impact of positive feedback is even more straightforward for these users. Kraut et al. (2011) claim that rewards motivate contributions and both gilds and upvotes can be seen as rewards.
@@
Regardless of whether our treated users were intrinsically or extrinsically motivated, positive feedback enhances that motivation (Kraut et al., 2011). Our findings confirm that users who contribute posts that get highly-upvoted increase their posting rate, indicating a potential increase in motivation to participate. Consequently, we assert that positive feedback is crucial to platform success given its significant role in supporting user motivation.
@@
The sorting algorithms of feeds on platforms like Reddit are often opaque to users and referred to as “black boxes” since platforms do not disclose the metrics they use to generate these feeds. However, users may have hypotheses about how the sorting works (Eslami et al., 2015). For example, surveyed moderators reported providing various forms of positive feedback to boost the visibility of content, implying that signals like score and gilds affect a post’s position in an algorithmically-curated feed (Lambert et al., 2024). However, our research has shown that positive feedback mechanisms can go beyond influencing the sorting algorithm and actually impact the quality of content generated by users in the future.
Because gilds and upvotes both encourage users to contribute posts that receive higher score and more gilds, we conclude that both forms of positive feedback encourage users to contribute posts that will be better received by the community. This indicates that the community overall is more satisfied with contributions made by recipients of positive feedback. This reveals revealing the power communities already have to proactively encourage a better pool of content through positive reinforcement. Thus, though moderators have additional weight and power to make decisions and enforce norms in a community, they are not the only ones capable of effecting positive change.
Though this work focuses on the capacity of community-members to facilitate distributed moderation through existing interface signals, centralized moderation teams on Reddit can learn from our analysis to inform how they moderate. Prior work found that YouTube comments receiving a “heart” from the creator of a video resulted in increased positive feedback from the community as a whole (Choi et al., 2024). Thus, when moderators demonstrate their approval of a contribution through highlighting, the community at large may reward it with positive feedback mechanisms that we found to positively impact user-level outcomes.
@@
Moreover, members of centralized moderation teams can curate posts to be rewarded through distributed moderation. Since Reddit does not explicitly provide moderator-specific feedback mechanisms, we suggest moderation teams focus on influencing the visibility of posts which can impact the amount of positive feedback they receive. More specifically, moderators can pin high-quality posts or periodically consolidate posts in a moderator-curated list in the subreddit sidebar, both of which moderators have reported doing (Lambert et al., 2024).

### v2

@@ 4.1.1. Findings
The results of our DiD analysis are reported in Table 1. The increase column is calculated from the coefficient on the $a_{t}*\delta(d\geq 0)$ term which is the interaction between time and treatment group. This indicates the effect of treatment on treated users in the observation window compared to if they had not received the treatment, i.e., the counterfactual. To reduce the likelihood of spurious significant relationships, we perform a Bonferroni correction by multiplying the $p$-values of each coefficient by the product of the number of regression tests (8) and the number of independent variables per regression (4). To capture the effect size, we also report the Cohen’s $d$ measured between the difference in each dependent variable before and after treatment for each treated user compared to the same difference before and after placebo for each control user.
We find that both forms of treatment lead to large positive increases in treated users’ score on future posts compared to if they had not received the treatment. More specifically, receiving gold and being highly upvoted correlate with a user receiving 19.28% and 57.11% higher score than they would have, respectively. Both treatments have much more modest, but still significant, effects on the average number of gold received on future posts. The gold treatment encourages 1.63% more gold in the future while the score treatment encourages an increase of less than 1% more gold.
Table 1. Difference-in-differences regression analysis results. We report the percent increase observed in the dependent variable for each treated user compared to if they had never received the treatment. Regressions with Bonferroni-corrected $p$-values are reported with asterisks (* $p<0.05$, ** $p<0.01$, *** $p<0.001$). Cohen’s $d$ reports the effect size of treatment on the dependent variables. We see significant increases in score and gold as a result of either treatment, and both increases in text difference and daily post rate and a decrease in daily removal rate from the score treatment.
@@
Gold Avg. Score 19.28% *** 0.16
Avg. # Gold 1.63% *** 0.27
Avg. Text Difference -0.23% -0.05
@@
Interestingly, users treated with score post content 0.34% more distinct from their treated post than they would have if they had not received the treatment. They also post 2.24% more per day than they would have and experience a negative effect on removal rate. This demonstrates that being highly upvoted might encourage users to post more frequent positive contributions while learning high-level community norms and thus violating them less often.
@@
Overall, receiving positive feedback of either kind encourages users to contribute posts that are better-received by their communities, and in some cases, also promotes user engagement and norm acquisition. Given the increase in future positive feedback and the decrease in norm-violations, we say that being highly upvoted encourages users to post higher-quality content according to both the community and the centralized moderation team.
@@ 4.2. RQ1b: How long do the effects of positive feedback on user behavior last?
Figure 5. Average treatment effects (ATE) of each treatment on the five outcomes. The rightmost $y$-axes correspond to the ATE of the score outcome which has a much larger range. The “x” markers indicate the days at which each outcome measure reaches saturation (i.e., decreases or stays the same for two consecutive days). The figure demonstrates the lasting effect of the score treatment on each outcome compared to the gold treatment, where the ATEs mostly return to zero within two weeks.
@@
A plot with two subplots. Both plots show the average treatment effects for each outcome from the start of the baseline period to the end of the observation window. The leftmost plot shows these values for the gold treatment and the rightmost plot visualizes the daily ATE values for the score treatment. The ’x’ icons on the leftmost plot show that the ATE for daily removal rate reaches saturation first, followed by daily post rate, average score, and finally average number of gold. Text difference does not rebound for the gold treatment. In the rightmost plot, removal rate reaches saturation first, then average score, daily post rate, text difference, and average number of gold.
To determine how long the effect of each treatment lasts (RQ1b), we compute the Average Treatment Effect (ATE) for each of our five outcomes on each day in our study period. ATE is computed as the difference between the treated users’ change in outcome and the control users’ change in outcome before and after treatment. We visualize the ATE values in Figure 5 averaged across each day in the baseline and observation windows. For each outcome, we also identify the day at which it reaches saturation, meaning the ATE for the outcome stops increasing for two consecutive days after treatment. Formally, our saturation point is defined as the day $d$ such that $ATE(d)\leq ATE(d-1)\leq ATE(d-2)$. For the removal rate outcome, we reverse the definition of saturation point since we are interested in seeing decreases in norm violations. In other words, the saturation point is the first day $d$ for which $ATE(d)\geq ATE(d-1)\geq ATE(d-2)$. These points are denoted by red “x” symbols in Figure 5.
We observe that for the gold treatment, the ATE values for all five outcomes return to nearly zero by the end of the observation window, indicating minimal differences between treated and control users. For the users treated with score, however, the ATE of their score and post rate in the observation window remain well above zero at the end of the two-week observation window, even though both measures reach their saturation point by day 5. This indicates that the effect of being highly-upvoted on future upvotes and engagement may last longer than the effect of being gilded. The elevated scores in the observation window may also serve as additional treatments, creating a snowball effect.
@@ 4.2.1. Rebound
We notice that most treated users do not rebound on the score outcome within 14 days, reflecting the trend shown in Figure 5 that score is elevated throughout the entire observation window. Daily post rate follows this same trend regardless of the treatment. However, for the average number of gold, nearly all users in both treatments rebound to their baseline values within two weeks. This is understandable given the rarity of gilding in the dataset.
Finally, we observe that the percentage of users who rebound to baseline text difference and removal rates after either treatment is much more evenly split. This suggests that the treatments have a lasting impact on the content of some users’ future posts and teach some users about the norms more effectively than others. Our subsequent RQ2 analysis explores this trend in more depth.
Table 2. Summary table describing the number of treated users that rebound. By definition, users must post at least twice in the 2-week post-treatment window in order to rebound, which is true for 2,599 (60.2%) users treated with gold and 90,993 (35.27%) users treated with score. The table shows that users almost always rebound on their average number of gold, rarely rebound on their average score and daily post rate, and sometimes rebound on their text difference and daily removal rate.
@@
Gold Avg. Score 74 (1.72%) 2,525 (58.53%)
Avg. # Gold 2,547 (59.04%) 52 (1.21%)
Avg. Text Difference 1,768 (40.98%) 831 (19.26%)
@@
5. RQ2: How do the effects of positive feedback vary across different types of recipients?
@@ 5.1. RQ2a: What types of users experience stronger effects of receiving positive feedback?
We compute Individual Treatment Effects (ITE) of each of our treatments on all five outcomes to evaluate user-level differential effects. ITE is used in prior work to compute how much a single treated user is affected by the treatment (Saha et al., 2019b). In our work specifically, ITE is computed as the difference between a treated user’s change in behavior after treatment and the average change in all control users’ behavior from the same stratum after placebo. More precisely, for an individual treated user $u_{t,s}$ from stratum $s$, we compute the ITE for each outcome as follows:
$\textit{ITE}_{\text{outcome}}(u_{t,s})=$ $\Delta u_{t,s}(\text{outcome})$
$-\text{mean}(\Delta u_{c,s}(\text{outcome}):u_{c,s}\in\mathcal{D}_{\textit{control,}s})$
@@
We then perform one linear regression analysis for each treatment-outcome pair using the outcome’s ITE as the dependent variable. Baseline values for our five outcomes are independent variables along with the author’s age, karma (i.e., the sum of scores on all posts) in the month prior to the baseline window, stratum, and the strength of treatment (i.e., number of gold or score on treated post). We take the log of each numerical independent variable.
@@ 5.1.1. Findings
Table 3 reports the coefficients from this regression. The bold coefficients are those we discuss in this section. All significance tests are Bonferroni-corrected based on the number of independent variables and regressions run. For the gold treatment, each $p$-value is multiplied by 40 (5 regressions, 8 IVs) and for score, each $p$-value is multiplied by 45 (5 regressions, 9 IVs). This regression analysis demonstrates that the strength of either treatment may lead to stronger effects. For example, posts treated with more gold are positively correlated with the treated user experiencing stronger effects on their future number of gold. Similarly, users who are treated with higher score have more extreme effects on their future positive feedback than users with more modest treatments. With this score treatment, however, we see that being treated with lower score also correlates with a larger effect on removal rate.
We also see trends aligned with the ideas of satiation and deprivation introduced by Miltenberger (2016). Specifically, our regression shows that users who have higher baseline average score experienced diminished effects of either treatment on their future score. Similarly, users with a higher average number of gold in their baseline window see less dramatic effects of the score treatment on their future number of gold. Both these trends reflect the idea that users are more strongly impacted by stimuli (e.g., gold or upvotes) if they have been deprived of that stimulus.
@@
Younger recipients of the score treatment also experience smaller increases in text difference, meaning newcomers post content in the future more similar to the post that received the reward compared to more established Reddit users.
Table 3. Individual Treatment Effects (ITE) regression results reported with Bonferroni-corrected $p$-values (* $p<0.05$, ** $p<0.01$, *** $p<0.001$). Coefficients with $p>0.05$ are excluded. Cells that are grayed out indicate independent variables not present in the regression. Bold coefficients are discussed in Section 5.1.1.
Treatment $\rightarrow$ Gold Score
IVs $\downarrow$ ITE $\rightarrow$ Score # Gold Text Diff. Post Rate Rem. Rate Score # Gold Text Diff. Post Rate Rem. Rate
Author Age 4.6E-04 *** -9.7E-03 *** -2.6E-03 ***
@@
Baseline Text Diff. -5.5E-01 *** 2.0E-01 ** -1.5E-01 *** -5.9E-03 *** -8.0E-01 *** -3.2E-02 *** 2.5E-01 ***
Baseline Post Rate -2.3E-02 *** -8.8E-01 *** -4.2E-03 *** -7.2E-03 *** -2.0E-01 *** 3.2E-01 ***
Baseline Rem. Rate -7.1E-02 * -2.6E-01 *** 2.2E+00 *** 4.4E-02 *** -3.1E-01 *** -2.4E+00 ***
Karma Last Month 1.1E-01 *** 1.8E-02 *** 5.4E-04 *** 1.0E-03 *
Gold Received 3.0E-02 **
Score Received 9.7E-02 *** 5.7E-04 *** -3.0E-03 ***
Baseline Avg. # Gold -5.0E-01 ***
R-squared 0.111 0.052 0.395 0.109 0.080 0.108 0.183 0.579 0.058 0.250
@@
Regression results for the ten regressions measuring the effect of the independent variables on the ITE values for each treatment-outcome pair.
@@ 5.2. RQ2b: What types of communities experience stronger average effects of receiving positive feedback?
$\textit{CTE}_{\text{outcome}}(r)=$ $\text{mean}(\Delta u_{t,r}(\text{outcome}):u_{t,r}\in\mathcal{D}_{\textit{treated,}r})$
$-\text{mean}(\Delta u_{c,r}(\text{outcome}):u_{c,r}\in\mathcal{D}_{\textit{control,}r})$
@@
Similar to the previous ITE analysis, we use our CTE values as dependent variables for ten linear regression analyses with subreddit-level monthly activity statistics from our entire study period as independent variables. This includes post rate, average number of gold, average score, the number of unique authors, and the number of subscribers.
@@ 5.2.1. Findings
After Bonferroni correction, these regressions yielded only two significant coefficients revealing that subreddits that gild more often see larger effects of the score treatment on their future number of gold and text difference. The lack of significant results indicates that the treatments have similar effects regardless of the community in which they occur.
@@ 6. Discussion
Our work contributes a computational approach and a causal inference framework through which researchers can examine the differential effects of treatment outcomes. We outline a method of identifying treatment and control posts, performing a difference-in-differences analysis, and determining user- and community-level factors that influence the strength of the treatment effect. This methodology can be adapted with different treatments, outcomes, and platforms to fit researchers’ needs. Additionally, while we focused on two types of positive feedback available on Reddit, upvotes in particular are almost universally applicable across social media platforms and have analogous features on Facebook, Instagram, X, Stack Overflow, and YouTube, among others. This ubiquity of upvotes enhances the generalizability of our work. Researchers interested in other platforms can use these findings to inform research into analogous features, and other forms of formal feedback, and extend this work with analyses of positive feedback signals not available on Reddit.
@@
6.1. Prioritize highlighting contributions from users recently deprived of positive feedback.
A key differential effect we observed relates to the principle of deprivation in positive reinforcement theory. The principle states that a stimulus has larger impacts on subjects deprived of it (Miltenberger, 2016). Our findings confirm this theory in the context of positive feedback on Reddit, highlighting that users who had not recently received gold experienced stronger effects of being gilded on their future gold received. Similarly, users who had not recently been highly-upvoted experienced larger increases in their future score when experiencing the score treatment.
@@
We recommend that designers and HCI researchers consider the principles of satiation and deprivation to best utilize valuable community resources like time and effort. Specifically, we suggest that they enable moderators and community members to maximize the potential effects of their positive feedback by building on work highlighting users’ history of toxicity (Im et al., 2020) and instead providing signals of how deprived or satiated an author is, or by prioritizing content in social media feeds contributed by authors deprived of positive feedback.
6.2. Provide mechanisms to bring attention to high-quality newcomer contributions.
@@
We showed that positive feedback in the form of upvotes results in the recipient posting fewer norm-violating posts than they would have without the treatment, implying that recipients are learning community norms. Additionally, newer accounts see stronger effects on their norm-adherence after treatment compared to more established Reddit users indicating that platform-level newcomers benefit most from positive feedback in the context of norm acquisition. Newcomers also more closely match the text of their future posts to the one that was rewarded, suggesting the power of positive feedback to shape the actual content of newcomer contributions.
@@
To enable this norm acquisition, we call on designers and HCI researchers to explore explicit ways for moderation teams to encourage established community-members to review newcomer contributions and provide positive feedback when applicable. For example, this may be a highlighting mechanism similar to YouTube creator hearts which have been shown to increase positive engagement for hearted comments (Choi et al., 2025). Designers may also consider building off a prior intervention that nudges users to preemptively reflect on their participation (Chang et al., 2022) and build a system to nudge users to specifically engage with newcomer content. Prior work on visualizing conversational metrics (Choi et al., 2023) can be extended to develop tools that guide moderator attention towards desirable behavior. Finally, platforms can incorporate a dedicated feed of newcomer contributions within communities to enable easy discovery of newcomer content.
@@
User motivation can be intrinsic or extrinsic (Kraut et al., 2011). In the context of our research, intrinsically-motivated users contribute high-quality content because they enjoy doing so. Kraut et al. (2011) highlight performance-based positive feedback as an especially effective method of enhancing users’ intrinsic motivation. Therefore, users who receive community-wide approval through positive feedback may experience increased intrinsic motivation to participate.
Extrinsically-motivated users may be more interested in making high-quality posts as a means to increase their reputation (i.e., karma) in the community. Rewards such as gold and upvotes are extrinsic motivators in themselves, which can help these types of users feel more motivated to participate (Kraut et al., 2011).
@@
Regardless of whether our treated users’ motivation was intrinsic or extrinsic, it was enhanced by positive feedback (Kraut et al., 2011). Our findings confirm that users who contribute posts that get highly-upvoted increase their posting rate, indicating a potential increase in motivation to participate. Consequently, we assert that positive feedback is crucial to platform success given its significant role in supporting user motivation.
@@
The sorting algorithms of feeds on platforms like Reddit are often opaque to users and referred to as “black boxes” since platforms do not disclose the metrics they use to generate these feeds (Chan et al., 2024). However, users may have hypotheses about how the sorting works (Eslami et al., 2015). For example, surveyed moderators reported providing various forms of positive feedback to boost the visibility of content, implying that signals like score and gold affect a post’s position in an algorithmically-curated feed (Lambert et al., 2024a). However, our research has shown that positive feedback mechanisms can go beyond influencing the sorting algorithm and actually impact the quality of content generated by users in the future.
Because gold and upvotes both encourage users to contribute posts that receive higher score and more gold, we conclude that both forms of positive feedback encourage users to contribute posts that will be better received by the community. This indicates that the community overall is more satisfied with contributions made by recipients of positive feedback. This reveals the power communities already have to proactively encourage a better pool of content through positive reinforcement. Thus, though moderators have additional weight and power to make decisions and enforce norms in a community, they are not the only ones capable of effecting positive change.
Though this work focuses on the capacity of community-members to facilitate distributed moderation through existing interface signals, centralized moderation teams on Reddit can learn from our analysis to inform how they moderate. Prior work found that YouTube comments receiving a “heart” from the creator of a video resulted in increased positive feedback from the community as a whole (Choi et al., 2025). Thus, when moderators demonstrate their approval of a contribution through highlighting, the community at large may reward it with positive feedback mechanisms that we found to positively impact user-level outcomes.
@@
Moreover, members of centralized moderation teams can curate posts to be rewarded through distributed moderation. Since Reddit does not explicitly provide moderator-specific feedback mechanisms, we suggest moderation teams focus on influencing the visibility of posts which can impact the amount of positive feedback they receive. Moderators can pin high-quality posts or periodically consolidate posts in a moderator-curated list in the subreddit sidebar, both of which moderators have reported doing (Lambert et al., 2024a).